OpenStudy (anonymous):

Solve for x 4x+12)º+(7x+3)º+6yº = 180º

OpenStudy (anonymous):

\[(4x+12)º+(7x+3)º+6yº = 180º\] I think that in place of 6yº it should be 6xº

OpenStudy (anonymous):

(4x+12)º+(7x+3)º+6xº=180º 17xº +15º =180º i.e. 17xº =180º -15º i.e. 17xº =165º i.e. x =165º/17 i.e. x= 10º i.e. x= 9.7058º i.e. x= 9.71º